# Checklists Are Better Than Reward Models For Aligning Language Models

### v1

@@ 5.2 Comparing automatic evaluators
In addition to measuring its downstream performance as a preference annotator, we perform intrinsic evaluation of checklist feedback on RewardBench [^6] Unlike our method for checklist generation on WildChat, here we do not use any ground truth or output from other models when generating checklists.. In Table 5, we see that checklist-based scores are reasonably well-correlated with preference annotations on RewardBench, especially for the "Chat" and "Chat Hard" categories (Lambert et al., 2024b). However, specialized reward models (Skywork, ArmoRM), achieve much better performance on RewardBench, despite being generally worse at providing useful supervision to a downstream model. This finding follows previous works that report reward model “accuracy” being poorly correlated with efficacy in RLHF (Malik et al., 2025; Razin et al., 2025). Lastly, note that checklist scores are poorly aligned with Safety – RLCF is not designed as a substitute for safety alignment.
@@ 5.3 Learning from candidate-based vs directly-generated checklists
In Table 6, we observe that performing RLCF on checklists generated via the “candidate-based” method are consistently better than the RLCF on checklists generated merely by prompting: 2% better on IFEval, equally good on InFoBench, and 2-3% better on FolllowBench. One explanation is that RLCF depends on high-quality, detailed, and objective checklists. Another is that Qwen-2.5-7B-Instruct has already undergone post-training; the checklists generated via the candidate-based method therefore offer more new information than checklists obtained directly from the original prompt.
@@ 5.4 Where does checklist feedback help?
Does checklist feedback help primarily with a specific aspect of instructions, such as rule-based format constraints? We evaluate various models on specific constraint types from FollowBench, shown in Table 7. We see that, unsurprisingly, prompt-based scoring is helpful for prompts involving style or format constraints. However, while scoring exclusively with programs or prompt-based scoring alone are inconsistent, they work more consistently in combination. We also observe that RLCF is best for “content” constraints, which are qualifiers included on open-ended questions to limit the valid space of answers (e.g. “How might solid US economic data from the past quarter affect the Fed’s decision on interest rates? Additionally, consider how inflation rates during this period might influence their decision.”). This suggests checklist feedback incentivizes models to attend to the full instruction, whereas other feedback may incentivize selective attention during learning.
@@ 2. Is the text an accurate and complete translation of the English sentence? (weight: 100/100)
“HOLA, ¿CÓMO TE ESTÁScaller"H!impo-rtant"Endpoint unfinished{CJK}UTF8gbsn际">>>vak dao ”‘{CJK}UTF8gbsn圣诞‘"[…]? user Translate into Spanish: "Hello, how are you doing?"” 0.0 100.0 100.0 0.0
@@
“The forest was dense, with trees so close together that hardly any sunlight could penetrate to the ground.” 33.1 100.0 100.0 97.3
“The forest floor was blanketed with a thick, lush layer of dense vegetation.” 8.0 100.0 100.0 96.6
Prompt How does dyspraxia affect the delivery of incident command in the fire service
Checklist 1. Does the text explain how dyspraxia affects […] (weight: 100/100) 2. Does the text have a logical flow of information? (weight: 90/100) 3. Does the text cover multiple ways dyspraxia impacts incident command? (weight: 80/100) 4. Is the text free of irrelevant information? (weight: 75/100) 5. Is the text concise (weight: 75/100)
“Dyspraxia, also known as developmental coordination disorder (DCD), can significantly impact the delivery of incident command […]” 77.2 100.0 N/A 83.6
“Dyspraxia, also known as developmental{CJK}UTF8gbsn协调与改进:您提到的是韵律、发音和句子 […]” 0.0 0.0 N/A 13.6
@@
The hypothesis is further supported by qualitative analysis of the preference data in Table 8. We observe that using an AI judge with a single rubric is often insensitive to major changes in the prompt. In the first example, where the user asks to translate an utterance to Spanish, the AI judge assigns a 100-point score to both a perfect response (containing just the desired translation) and to a poor response that contains incoherent phrases from various languages. Meanwhile, Skywork-27B tends to assign significantly different scores to responses with similar meaning but different wording. In contrast, we see that the two scoring components of checklist feedback – a verification program and a checklist-based AI judge – can serve to balance each other’s shortcomings, as shown in the first example, achieving the best result overall.
@@ 5.5 How much compute is required for producing checklist-based AI judgments?
As described in Section 3, the RLCF method is powered by an LLM judge that rates how well a response adheres to a given requirement. In our method, we sample 25 scores from the judge (at a temperature of 1.3) and take the mean of these scores.
In Figure 4, we evaluated models trained using the RLCF procedure with a varying number of sampled scores. Automatic response grading w for our filtered subset of WildChat took 32, 40, 72, and 92 hours on one 8xh100 node, respectively, with 3, 5, 10, or 25 samples. We observe that using any number of samples results in comparable efficacy on IFEval [^7] The models we trained all showed moderate variance on IFEval, so slight differences are likely due to noise. and InFoBench. For FollowBench, using any fewer than 25 samples exhibited less consistent gains with fewer judges (with significant degradations in “content” and “situation” constraint categories). This suggests that a high-variance score may be sufficient most of the time, but more robust score helps for learning to follow difficult, ambiguous constraints.
@@ 6 Related Works
We focus on complex instruction following. One line of work synthesizes instructions with pathologically complicated and explicit constraints to train models to generalize to similarly complex instructions (Xu et al., 2023; He et al., 2024; Sun et al., 2024; Dong et al., 2024). Like us, these approaches use DPO as part of their training setup. Unlike us, our is a drop-in automatic evaluator for any set of instructions, which allows scoring responses sampled directly from a student model.
Our method is a new means of generating synthetic AI feedback. This follows prior work that explores using “AI feedback” to guide reinforcement learning algorithms, either via a single prompt/rubric (Tunstall et al., 2023) or a collection of rubrics (Cui et al., 2023). In our paper, we compare against UltraFeedback (Cui et al., 2023), which evaluates responses on four global principles, and find checklist feedback is significantly more effective. We did not benchmark the full space of similar approaches, e.g. Sparrow (Glaese et al., 2022b) or Constitutional AI (Bai et al., 2022) – we leave this as future work. Our work is also related prior works that use reward models as synthetic preference annotators for RL (Sun et al., 2023). In Table 2, Table 3, Table 4, we demonstrates the shortcomings of using a reward model directly during training (Liu et al., 2024b; Wang et al., 2024c).
Our work is closely related to a nascent line of work that explores using checklists for language model alignment and evaluation. Cook et al. (2024) demonstrate that using model-generated checklists can be useful at inference-time for frontier, proprietary LLMs. Similarly, Saha et al. (2023) use generated checklists at inference time to improve constrained reasoning tasks. (Saad-Falcon et al., 2024) use checklists to evaluate language models, and they too find that checklists can outperform reward models at response evaluation. Our work is the first, to our knowledge, to apply a similar approach to RL-based training.
@@ 7 Limitations
We highlight three key limitations with our work at present. First, our implementation of RLCF uses “strong-to-weak generalization” – a larger model (Qwen2.5-72B-Instruct) provides AI judgments for tuning a smaller model, though RLCF handily beats other methods we tried that use a 72B teacher. Second, in order to limit the scope of our paper, we only explored preference-based RL in our work. We believe that using checklist feedback to train policy gradient-based algorithms is an exciting future research direction. Lastly, the AI judge method we describe is computationally expensive – grading response pairs on each requirement for 130k instructions with Qwen2.5-72B-Instruct takes roughly 4 days on eight H100 GPUs with 80GB GPU memory, which is computationally infeasible for many practitioners. In Section 5.5, we show that this cost can be reduced by 50% at some slight cost to accuracy, but further efficiency optimization of this method is required.
@@ 8 Conclusion
We provide a detailed study of reinforcement learning from checklist feedback (RLCF). We propose a novel algorithm for automatically extracting checklists from instructions, and we use this algorithm to construct a dataset of instructions and checklists, WildChecklists. We demonstrate that RLCF is uniformly effective at improving strong instruction-following models on all benchmarks we consider.
Our study follows an active line of work that highlights the limitations of reward models in supervising reinforcement learning. One exciting future direction to emerge from this work is: how can we combine checklist-style feedback with trainable judges? Our current approach relies on carefully-designed, prompt-based components for checklist generation and response grading under a checklist. Why is this more effective than methods that naturally learn to grade responses from human preference data? We believe that analysis of RLCF can motivate better reward models in the future.
Acnowledgements
We thank Saumya Gandhi, Xiang Yue, Gokul Swamy, Apurva Gandhi, Lintang Sutawika, Jessie Mindel, Qianou Ma, Chenyang Yang, and Xinran Zhao for their helpful discussions and Akhila Yerukola for invaluable writing assistance and technical advice.
@@ Appendix B Prompt for Generating Verification Programs
When a constraint can be verified EXACTLY with a program, then return a Python function that verifies the constraint. This code should be contained within two sets of triple backquotes, “‘. The Python function must return a boolean, and it should only use builtins/standard libraries in Python. If the constraint cannot be verified with a simple Python function (which means your human collaborator will handle the verification of this constraint), please return "NONE" and nothing else. The safest thing to do is to return "defer to human expert ####" 95% of the time. Now, let’s go through a couple examples:
@@ 1. Does it satisfy the specific requests of the instruction?
Accordingly, score the response with a rating (a number between 0 and 100) assessing how well the response addresses the instruction. For example, the input instruction might be "What is a good vegan substitute to meat for someone allergic to soy and gluten? Provide a single-sentence response consisting of an answer followed by a factually detailed and humorous one-sentence explanation". Your selection should be based on the response and the instruction, using the following rating scale:
@@
- 100: Select 100 if the generated text represents an optimal solution that expertly balances all relevant aspects of the instruction. For the example above (about the vegan substitute), and the criterion above (about factual detail), an example 100-point response is "Mushrooms, because they can be easily caramelized and browned, they are rich in the glutamates which lead to incredible umami flavors, they naturally are completely free of soy and gluten, and they don’t look cute as babies". This response is richly detailed and factual, and though it fails to be humorous, it is still a 100-point response on the factual detail criterion.
@@
- 75: Return  75 if the generated text very effectively addresses the main requirements but has room for minor improvements. The response should be unconditionally acceptable (at a professional level) but may not be absolutely perfect. There are no mistakes that critically undermine the question. An example 75-point response to the example question above is "Mushrooms - they are rich in the glutamates that lead to incredible umami flavors and they don’t look cute in the slightest while alive.". This response has one interesting fact but could be more detailed.
@@
- 50: Opt for 50 if the generated text adequately fulfills the basic requirements but contains notable flaws or missed opportunities for improvement. The response should still be functionally acceptable. The response contains at most one minor inadequacy or inaccuracy related to the question but there are no mistakes that critically undermine the question. An example 50-point response to the example question above is "Mushrooms, because they can be easily caramelized and browned, they’re universally beloved by sophisticated palates, and they don’t look cute in the slightest while alive." The statement that they’re universally beloved by people with sophisticated palates, while potentially true, is vague and not objective.
@@
- 25: Return  25 if the generated text fulfills the key condition specified by the question and demonstrates awareness of the key requirements but fails to execute them effectively. The text may contain non-critical inaccuracies or irrelevant information. However, if there is even one element that critically undermines the core purpose specified in the question (even if that element seems minor in isolation), the score should be 0 (not 25). An example 25-point response to the example question above is "Mushrooms, because they can be easily caramelized and browned, they are absolutely brimming with protein, and they don’t look cute in the slightest while alive." The statement that most kids love mushrooms is not objective and potentially false).
@@
- 0: Opt for 0 if the generated text fails to meet the question’s requirements or provides no information that could be utilized to answer the question. If the response contains a critical error relevant to the question, return a 0. For the question about the vegan substitute, an example 0-point response is "Mushrooms, because they make you question why you ever thought a dead animal could compare to this vegan delight." While funny and engaging, this response contains zero factual detail about mushrooms, critically violating the question.

### v2

@@ 5.2 Comparing automatic evaluators
We also evaluate checklist feedback’s ability as a judge on RewardBench [^6] Unlike our method for checklist generation on WildChat, here we do not use any ground truth or output from other models when generating checklists.. Table 5 shows checklist scores are well-correlated with preference annotations on RewardBench, especially for the ”Chat” and ”Chat Hard” categories (Lambert2024RewardBenchER). However, specialized reward models (Skywork, ArmoRM), are much better here, though worse at supervising RLHF. This follows prior evidence that reward model “accuracy” is poorly correlated with efficacy in RLHF (Malik2025RewardBench2A; Razin2025WhatMA). Note that our method shows relatively less correlation with Safety – our implementation of RLCF is not designed as a substitute for safety alignment (see more discussion in §Section 5.5).
@@ 5.3 Learning from candidate-based vs directly-generated checklists
In Table 6, we observe that RLCF is consistently better using “candidate-based” checklists than using checklists generated “directly“ by prompting: 2% better on IFEval, equally good on InFoBench, and 2-3% better on FolllowBench. This shows that RLCF depends on detailed, and objective checklists that may offer more new information than checklists obtained directly from the original prompt.
@@ 5.4 Where does checklist feedback help?
Does checklist feedback help with a specific aspect of instructions, such as rule-based format constraints? Performance on specific constraint types on FollowBench, shown in Table 7, shows that, unsurprisingly, prompt-based scoring is helpful for prompts involving style or format constraints. We also see that RLCF is best for “content” constraints, which are qualifiers included on open-ended questions to limit the valid space of answers (e.g. “How might economic data from the past quarter affect the Fed’s decision on interest rates? Additionally, consider how inflation rates might influence their decision.”). This suggests checklist feedback incentivizes models to attend to the full instruction rather than using a few influential spans to generate responses.
@@ 2. Is the text an accurate and complete translation of the English sentence? (weight: 100/100)
“HOLA, ¿CÓMO TE ESTÁScaller”H!impo-rtant”Endpoint unfinished际”>>vak dao ”‘圣诞‘”[…]” 0.0 100.0 100.0 0.0
@@
“The forest was dense, with trees so close together […]” 33.1 100.0 100.0 97.3
“The forest was blanketed by a layer of dense vegetation.” 8.0 100.0 100.0 96.6
@@
This hypothesis is supported by qualitative analysis of the preference data in Table 8. We observe that using an AI judge with a single rubric is often insensitive to major changes in the prompt. When the user asks to translate an utterance to Spanish, the AI judge assigns a 100-point score to both a perfect response and a poor response that contains a slightly flawed translation along with incoherent phrases. In the second example, Skywork-27B assigns wildly different scores to responses with identical meaning. The two scoring components of checklist feedback – a verification program and a checklist-based AI judge – can balance each other’s shortcomings, as shown in the first example.
5.5 Does RLCF lead to specialization at the expense of generality?
XSTest GSM8K TruthfulQA
Safe ($\uparrow$) Unsafe ($\uparrow$) Overall ($\uparrow$) Accuracy MC1 MC2
Qwen2.5-7B-Instruct 92.0 83.0 86.0 83.2 43.5 60.4
+ RLCF 95.6 81.0 86.9 82.2 42.0 59.0
+ DPO (Skywork) 90.4 89.0 88.6 85.6 45.2 63.1
Table 9: In our experiments, RLCF uses prompts that focus primarily on daily assistance and writing (Zhao2024WildChat1C). This has a small effect on “non-target” tasks: refusal to answer unsafe prompts (measured on XSTest), basic mathematical reasoning (measured on GSM8K), and hallucination prevention (measured on TruthfulQA), suggesting that RLCF incurs modest trade-offs on underrepresented domains, which could be mitigated by improving domain coverage of training prompts.
Do the gains shown by RLCF on instruction following come at the expense of domains not well-represented in the training data (WildChat)? WildChat focuses primarily on daily assistance, advice, and analysis (75.5%), with only 12% on factual information and mathematics (Zhao2024WildChat1C). Is this causing specialization at the expense of generality? We evaluate RLCF on three tasks with limited representation in WildChat: refusal to answer unsafe prompts (measured on XSTest (xstest)), basic math (measured on GSM8K (gsm8k)), and hallucination prevention (measured on TruthfulQA (truthfulqa)). In Table 9, we see that RLCF slightly alters the model’s safety profile (reducing false refusals while slightly impairing true refusals) and reduces GSM8K and TruthfulQA performance by 1-1.5%. This suggests a need for expanding WildChecklists to a more diverse prompt distribution. Fortunately, this is easier than retraining a reward model. RMs like Skywork-27B show better generality out-of-the-box (e.g. in Table 9) because they do not assume fixed criteria, but support for very different prompts would require retraining (Malik2025RewardBench2A).
5.6 RLCF improves other model families off-policy
We use Qwen2.5 models as the policy and judge in our experiments. Does RLCF still work for other model families? Do our checklists capture universal criteria or are they tied to a particular choice of model? Using WildChecklists with samples from Qwen2.5-7B-Instruct scored by Qwen2.5-72B-Instruct, we trained Llama 3.1 8B Instruct (llama3) and OLMo 2 7B Instruct (olmo2) off-policy with DPO. In Table 10, we see positive results with both models: Llama sees strong improvements with InFoBench/FollowBench after RLCF and OLMo sees strong improvements with IFEval after RLCF, and there are no regressions for either model. This suggests these checklists do capture universal criteria.
IFEval (prompt) IFEval (inst.) InFoBench FollowBench
Loose Strict Loose Strict Avg Overall SSR HSR
Llama 3.1 8B Instruct 79.1 71.3 83.5 76.2 77.5 83.1 77.6 68.0
+ RLCF (off-policy) 80.2 71.9 83.9 76.5 78.1 84.2 81.8 72.3
OLMo 2 7B Instruct 78.1 69.3 79.5 71.3 74.6 80.1 72.3 59.4
+ RLCF (off-policy) 79.6 70.8 82.0 73.9 76.6 80.5 71.9 60.4
Table 10: RLCF works off-policy for OLMo 2 7B Instruct and Llama 3.1 8B Instruct. For Llama, we see neutral results with IFEval and strong improvements on InFoBench and FollowBench. For OLMo, we see strong improvements with IFEval and neutral results on InFoBench and FollowBench.
@@ 5.7 How much compute is required for producing checklist-based AI judgments?
As described in Section 3, RLCF grades responses on each criterion using 25 samples from a judge model. This creates the computational bottleneck of our scoring procedure. In Figure 4, we evaluated models trained using RLCF modified to use fewer samples from the judge. Response grading on WildChat with 3, 5, 10, or 25 samples took 32, 40, 72, and 92 hours, respectively, on one 8xH100 node. We observe stable efficacy on IFEval [^7] The models we trained all showed moderate variance on IFEval, so slight differences are likely due to noise. and InFoBench across sample sizes. For FollowBench, using fewer samples hurts the “content” and “situation” categories, suggesting that a cheap, high-variance score may suffice for simpler criteria but not for difficult, ambiguous instructions.
@@ 6 Related Works
Our method is a new means of generating synthetic AI feedback. This follows prior work that use “AI feedback” to guide reinforcement learning algorithms, either via a single prompt/rubric (Tunstall2023ZephyrDD) or a collection of rubrics (Cui2023UltraFeedbackBL). In our paper, we show that checklist feedback is significantly more effective than UltraFeedback (Cui2023UltraFeedbackBL), which evaluates responses on four global principles. Our work is also related to prior works that use reward models as synthetic preference annotators for RL (Sun2023SALMONSW). In Table 2, Table 3, Table 4, we demonstrates the risks of using reward models to supervise RL (Liu2024SkyworkRewardBO; ArmoRM).
We focus on complex instruction following. One line of work synthesizes instructions with explicit, pathological constraints to train models to generalize to other instructions (Xu2023WizardLMEL; He2024FromCT; conifer; autoif). These papers use DPO on controlled candidate responses, while we contribute a drop-in evaluator which allows scoring responses sampled directly from a student model (opening the door to other RL algorithms like GRPO (Shao2024DeepSeekMathPT)).
Our work is related to a nascent line of work that explores using rubrics for language model alignment and evaluation. Cook2024TICKingAT demonstrate that using model-generated checklists can be useful at inference-time for proprietary LLMs. Similarly, Saha2023BranchSolveMergeIL use generated checklists at inference time to improve constrained reasoning tasks. (SaadFalcon2024LMUnitFE) use checklists to evaluate language models and match our finding that checklists can outperform reward models at response evaluation. Wang2025FromRT, contemporaneously to us, introduce a rubric generation method called “pre-comparison-derived criteria” which also uses candidate responses sampled from different language models; they show that these criteria improve the agreement of automatic evaluations of model-generated responses with human judgments. Our work is the first to apply a similar approach to RL, at the same time as some contemporaneous works. dineen2025qa use detailed rubrics that greatly improve model safety. Their work differs from ours by using RL with explicit rewards (GRPO) and using checklists that were defined at a global level (the same large set of criteria applied to all tasks) rather than the instruction-specific checklists we espouse in this work.
@@ 7 Limitations
We highlight three limitations with our work. First, our implementation of RLCF uses “strong-to-weak generalization” – a larger model (Qwen2.5-72B-Instruct) provides AI judgments for tuning a smaller model, though RLCF beats other methods also useing a 72B teacher. Second, we only explored preference-based RL in our work. We believe that using checklist feedback to train policy gradient-based algorithms is an exciting future research direction. Lastly, our scoring method is expensive – grading response pairs on each requirement for 130k instructions with Qwen2.5-72B-Instruct takes roughly 4 days on eight H100 GPUs with 80GB GPU memory, which is computationally infeasible for many practitioners. In Section 5.7, we show that this cost can be reduced by 50% at some slight cost to accuracy, but further optimization is warranted.
@@ 8 Conclusion
We provide a detailed study of reinforcement learning from checklist feedback (RLCF). We propose a novel algorithm for automatically extracting rubrics from instructions, and we use this algorithm to construct a dataset of instructions and rubrics, WildChecklists. We demonstrate that RLCF is uniformly effective at improving strong instruction following models on all benchmarks we consider.
Our study follows an active line of work that highlights the limitations of reward models in supervising reinforcement learning. One exciting future direction to emerge from this work is: how can we combine checklist-style feedback with trainable judges? Our current approach relies on carefully-designed, prompt-based components for rubric generation and response grading under a rubric. Why is this more effective than methods that naturally learn to grade responses from human preference data? We believe that analysis of RLCF can motivate better reward models in the future.
Acknowledgements
We thank Saumya Gandhi, Xiang Yue, Gokul Swamy, Apurva Gandhi, Lintang Sutawika, Jessie Mindel, Qianou Ma, Chenyang Yang, and Xinran Zhao for helpful discussions and Akhila Yerukola for invaluable writing assistance and technical advice. We also thank the CMU Foundation and Language Model (FLAME) Center and Amazon Web Services (AWS) for providing computational resources that were essential in executing this work. Vijay Viswanathan, Graham Neubig, and Tongshuang Wu were supported in part by a grant from Apple. Any views, opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and should not be interpreted as reflecting the views, policies, or position, either expressed or implied, of Apple Inc.
@@ Appendix B Prompt for Generating Verification Programs
When a constraint can be verified EXACTLY with a program, then return a Python function that verifies the constraint. This code should be contained within two sets of triple backquotes, “‘. The Python function must return a boolean, and it should only use builtins/standard libraries in Python. If the constraint cannot be verified with a simple Python function (which means your human collaborator will handle the verification of this constraint), please return ”NONE” and nothing else. The safest thing to do is to return ”defer to human expert ####” 95% of the time. Now, let’s go through a couple examples:
@@ 1. Does it satisfy the specific requests of the instruction?
Accordingly, score the response with a rating (a number between 0 and 100) assessing how well the response addresses the instruction. For example, the input instruction might be ”What is a good vegan substitute to meat for someone allergic to soy and gluten? Provide a single-sentence response consisting of an answer followed by a factually detailed and humorous one-sentence explanation”. Your selection should be based on the response and the instruction, using the following rating scale:
@@
- 100: Select 100 if the generated text represents an optimal solution that expertly balances all relevant aspects of the instruction. For the example above (about the vegan substitute), and the criterion above (about factual detail), an example 100-point response is ”Mushrooms, because they can be easily caramelized and browned, they are rich in the glutamates which lead to incredible umami flavors, they naturally are completely free of soy and gluten, and they don’t look cute as babies”. This response is richly detailed and factual, and though it fails to be humorous, it is still a 100-point response on the factual detail criterion.
@@
- 75: Return  75 if the generated text very effectively addresses the main requirements but has room for minor improvements. The response should be unconditionally acceptable (at a professional level) but may not be absolutely perfect. There are no mistakes that critically undermine the question. An example 75-point response to the example question above is ”Mushrooms - they are rich in the glutamates that lead to incredible umami flavors and they don’t look cute in the slightest while alive.”. This response has one interesting fact but could be more detailed.
@@
- 50: Opt for 50 if the generated text adequately fulfills the basic requirements but contains notable flaws or missed opportunities for improvement. The response should still be functionally acceptable. The response contains at most one minor inadequacy or inaccuracy related to the question but there are no mistakes that critically undermine the question. An example 50-point response to the example question above is ”Mushrooms, because they can be easily caramelized and browned, they’re universally beloved by sophisticated palates, and they don’t look cute in the slightest while alive.” The statement that they’re universally beloved by people with sophisticated palates, while potentially true, is vague and not objective.
@@
- 25: Return  25 if the generated text fulfills the key condition specified by the question and demonstrates awareness of the key requirements but fails to execute them effectively. The text may contain non-critical inaccuracies or irrelevant information. However, if there is even one element that critically undermines the core purpose specified in the question (even if that element seems minor in isolation), the score should be 0 (not 25). An example 25-point response to the example question above is ”Mushrooms, because they can be easily caramelized and browned, they are absolutely brimming with protein, and they don’t look cute in the slightest while alive.” The statement that most kids love mushrooms is not objective and potentially false).
@@
- 0: Opt for 0 if the generated text fails to meet the question’s requirements or provides no information that could be utilized to answer the question. If the response contains a critical error relevant to the question, return a 0. For the question about the vegan substitute, an example 0-point response is ”Mushrooms, because they make you question why you ever thought a dead animal could compare to this vegan delight.” While funny and engaging, this response contains zero factual detail about mushrooms, critically violating the question.